A portfolio produces the following returns in years 1-5:

Year Return (%)
1 4
2 -5
3 4
4 8
5 -1

What is the Sharpe ratio of this portfolio, if the average risk free rate over the same time period was 3%?

Enter answer accurate to 2 decimal places.
Answer #1

Sharpe ratio = (Rp - Rf)/ SD(p)

Rp = Average return of the portfolio = 2 (using excel formula average(r1....r5)

SD(p) = 5.05 (using excel formula stdev.s(r1....r5))

Sharpe ratio = (2 - 3%)/5.05 = -0.198 = -0.20 (rounded off to two decimals)